OpenStudy (anonymous):

Francine made a rectangular prism out of 40 cubic centimeters of clay. If its length is twice its width and its height is 1 centimeter more than its length, what is its length in centimeters? [Volume of a rectangular prism = length × width × height] a)1 b)2 c)4 d)6 e)-7

OpenStudy (danjs):

V = 40 cm^3 L = 2*W H = L + 1

OpenStudy (danjs):

V = L * W * H = 40

OpenStudy (danjs):

H = L + 1 and L = 2W so H = 2W + 1

OpenStudy (danjs):

V = 40 = 2W * W * (2W+1) Solve for W and use H = 2*W + 1

OpenStudy (danjs):

L = 2W H = 2W + 1 W = W V = L x W x H

OpenStudy (anonymous):

okay, so do i just plugging in the numbers for w in this? V = 40 = 2W * W * (2W+1)\

OpenStudy (danjs):

yes, solve for what W is

OpenStudy (anonymous):

and if it equals 40 then thats the answer?

OpenStudy (danjs):

no, you can solve that for W, then use that number to find the Length L by using L = 2*W

OpenStudy (anonymous):

oh so i dont plug the answers into that to find w, i just solve it like a regular equation?

OpenStudy (danjs):

40 = 2W^2*(2W+1) you will find W = 2 so L = 2*W = 4

OpenStudy (danjs):

There is nothing to plug in, You have to use that equation to find what W is then use that W to find what the L is.

OpenStudy (anonymous):

okay, i think i got it is it l=4?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

thank you

OpenStudy (danjs):

yep

OpenStudy (danjs):

Just read the question once or twice slow to get a feel for it, Then take it once sentence at a time, and try to translate it into equations. "Length is twice the width" L = 2*W